OpenStudy (anonymous):

if sin a is a given amount, and sin b is a given amount, then what quadrant is a+b in?

OpenStudy (anonymous):

what is given amount?

OpenStudy (anonymous):

I think it wont be in a quadrant but rather the x axis. IDK tho lol.

OpenStudy (anonymous):

cos a = -4/5 cos b = 8/17 sina=-3/5 sinb=-15/17 a is in quadrant 3 b is in quadrant 4

OpenStudy (anonymous):

from the given quantities find out the values of both cos(a+b) and sin(a+b). Then, depending on the signs of the values you get, you can determine the quadrant in which (a+b) lies.

OpenStudy (anonymous):

thank you!